# Device-independent certification of desirable properties with a confidence interval

### v1

@@ IV.1.2 Dimension certification via negativity certification
As mentioned in Section II.2.1, the correlation $\vec{P}_{\text{CHSH}}$ is insufficient to demonstrate any nontrivial dimension bound. Next, let us consider, instead, a correlation $\vec{P}_{\text{CGLMP}}$ derived by local measuring the partially entangled two-qutrit state
|$\ket{\Psi}=\frac{1}{\sqrt{2+\zeta^{2}}}(\ket{00}+\zeta\ket{11}+\ket{22}),\quad\zeta=\frac{1}{2}(\sqrt{11}-\sqrt{3}).$ (39a)
@@
where $\varphi^{A}_{x}=\frac{x}{2}$, $\varphi^{B}_{y}=(-1)^{y}\frac{1}{4}$, and {|$\{\ket{j}\}$ is the set of computational basis states. It is known [70, 66] that this strategy gives the maximal CGLMP Bell-inequality violation of $\mathcal{S}_{\text{\tiny CGLMP3}}=1+\sqrt{\frac{11}{3}}\cong 2.9149$. Moreover, the negativity of |$\ket{\Psi}\!\bra{\Psi}$ | can be easily evaluated to give $\cong 0.983$.
As with the case of $\vec{P}_{\text{CHSH}}$, we again use the numerically simulated Bell test data to perform a hypothesis testing for the null hypothesis of 1, but now with $\mathcal{N}_{0}\in\bm{N}_{0}=\{0.5,0.51,\cdots,0.98\}$. For the PBR protocol, the computation proceeds in exactly the same way as described above [see the paragraph containing Section IV.1.1]. However, for the martingale-based protocol, since we do not have an explicit expression like that shown in Eq. 37 for the CGLMP Bell expression, we compute an upper bound on $B_{\mathcal{H}}$ for each given value of $\mathcal{N}_{0}\in\bm{N}_{0}$ according to:
@@
where the CGLMP Bell coefficients are defined in Eq. 11. Meanwhile, since $P_{xy}=\frac{1}{4}$ and $\beta^{ab}_{xy}\in\{-1,0,1\}$, we again have $b_{\pm}=\pm 4$ for $I_{\text{\tiny CGLMP3}}$.
Figure 4: Average amount of negativity certifiable from the data observed in a Bell test generating $\vec{P}_{\text{CGLMP}}$. For details on how the plot is generated, see the caption of Fig. 2, but bearing in mind that for the martingale-based protocol, we now use in Eq. 25 the $B_{\mathcal{H}}$ determined from Section IV.1.2.
@@
and numerically maximize their CGLMP Bell-inequality using the heuristic algorithm given in [93]. We denote the corresponding correlation by $\vec{P}_{\tilde{\zeta}}$, compute the corresponding asymptotic confidence-gain rate for both protocols, and plot the results in Fig. 5.
Figure 5: Asymptotic confidence-gain rate $G^{\text{(prot)}}$ based on the quantum correlation $\vec{P}_{\tilde{\zeta}}$ derived from maximizing the CGLMP Bell-inequality-violation of |$\ket{\Psi(\tilde{\zeta})}$, Eq. 41. Here, we consider $\tilde{\zeta}=\{0,0.02,...,1\}$ and the composite hypothesis 1, with $\mathcal{N}_{0}=0.5$. The gain rate for the martingale-based protocol is computed from Eq. 27 using the CGLMP Bell function and the bound $B_{\mathcal{H}}$ determined in Section IV.1.2 whereas that for the PBR protocol is evaluated from Eq. 36 and setting $\vec{P}_{\mathcal{Q}}$ to $\vec{P}_{\tilde{\zeta}}$. The dashed blue line corresponds to the value $\tilde{\zeta}=\zeta$, cf. Eq. 39a.
Interestingly, even though Fig. 4 suggests that the CGLMP Bell function is very effective in providing a good $p$-value bound against hypothesis 1,  Fig. 5 clearly show that, asymptotically, it is not optimal. The results shown in Fig. 5 further suggest that among the family of two-qutrit states given in Eq. 41, the qutrit signature of |$\ket{\Psi(\zeta)}$, cf. Eq. 39a, could even the most prominent, when it comes to its DI certification using these hypothesis-testing techniques.
@@ IV.2 Entanglement depth certification
It is easy to verify that $\vec{P}_{\text{GHZ}}$ leads to a violation of the Mermin Bell inequality,  Eqs. 12 and 14, giving the algebraic maximum of $\mathcal{S}_{\text{\tiny Mermin}}=4$. For our simulations, we generate $N_{\text{tot}}=10^{4}$ trials for each of the 30 Bell tests, assuming a uniform distribution $P_{xyz}=\frac{1}{4}$ over all measurement settings $x,y,z\in\{0,1\}$ that satisfy $\text{mod}(x+y+z,2)=1$. Then, we test the data against the following composite hypotheses.
Hypothesis 2.
[…]
Hypothesis 3.
$\mathcal{H}_{\text{2-prod}}$: In every experimental trial, the underlying state is 2-producible, i.e., having an entanglement depth of $2$ of less.
For the martingale-based method, we use Eq. 25 with the Mermin Bell expression of Eq. 12 and the bounds given in Eq. 14, i.e., $B_{\mathcal{H}}=2$ for hypothesis 2 and $B_{\mathcal{H}}=2\sqrt{2}$ for hypothesis 3. Since $\beta^{abc}_{xyz}\in\{-1,0,1\}$, we again have $b_{\pm}=\pm 4$. Note that separable states can only generate Bell-local correlations [5], cf. Eq. 2. Thus, for the PBR protocol with hypothesis 2, the optimizing distribution $P^{\star}_{k}(abc|xyz)$ for the $k$-iteration can be obtained by solving [cf. Eq. 29]
@@
where $\rho_{i}$, with $i=\{1,2,3\}$ are meant to represent, respectively, the constituent of $\rho$ that is separable with respect to the $A|BC$, $B|AC$, and $C|AB$ bipartitions. In evaluating Section IV.2, we use level $\ell=1$ of the hierarchy introduced in [16]. For both hypotheses, we then evaluate
$R(a,b,c,x,y,z)=\frac{f_{\text{reg}}^{(k)}(abc|xyz)}{P^{\star}(abc|xyz)}$ (45)
for the computation of the test statistic $t_{k}$. In computing these $t_{k}$, we consider a block size of $N_{\text{blk}}=50$.
@@
$\mathfrak{P}^{\text{(pbr)}}_{\mathcal{H}_{\text{Sep}}}=0.414857\,N-22.902259,\,\,N\in[10^{2},10^{4}],$ (46)
$\mathfrak{P}^{\text{(mart)}}_{\mathcal{H}_{\text{Sep}}}=0.415037\,N,\quad N\in[0,10^{4}]$ (47)
@@
$\mathfrak{P}^{\text{(pbr)}}_{\mathcal{H}_{\text{2-prod.}}}=0.228124\,N-15.437662,\,\,N\in[10^{2},10^{4}],$ (48)
$\mathfrak{P}^{\text{(mart)}}_{\mathcal{H}_{\text{2-prod.}}}=0.228447\,N,\quad N\in[0,10^{4}]$ (49)
for the 2-producible hypothesis $\mathcal{H}_{\text{2-prod.}}$ of 3. Consequently, even if we only run the Bell test using the strategy of Section IV.2 for 100 trials, there is already sufficient data to certify genuine tripartite-entanglement with a confidence of at least $1-10^{-6}$.
@@ IV.3 Fidelity certification
Hypothesis 4.
[…]
Then, for any given $\theta$ and $\mathcal{F}_{0}\geq\cos\theta$, to apply the PBR protocol, we solve the optimizing distribution $P^{\star}_{k}(ab|xy)$ for the $k$-iteration [cf. Eq. 29] via:
@@
where the left-hand side of the last inequality in Eq. 51b consists of some specific linear combination of entries of $\chi_{\ell}[\rho]$, see [86] for details. Then, as with negativity certification, we can evaluate the PBR used in the computation of $t_{k}$ by replacing $f_{\text{reg}}(ab|xy)$ and $P^{\star}(ab|xy)$, respectively, by $f_{\text{reg}}^{(k)}(ab|xy)$ and $P^{\star}_{k}(ab|xy)$ in Eq. 30. Again, we consider a block size of $N_{\text{blk}}=500$. As for the martingale-based protocol, we first solve
@@
Let us start with the self-test of a Bell state, corresponding to $\theta=\frac{\pi}{4}$ in Eq. 17a. In this case, we use the CHSH Bell function specified in Eq. 4 and consider $\mathcal{F}_{0}\in\bm{F}_{0}=\{0.5,0.51,\cdots,0.99\}$. For both protocols, by systematically evaluating the $p$-value bounds from the data for each of these $\mathcal{F}_{0}$’s, we determine a lower bound on $\mathcal{F}_{\theta=\frac{\pi}{4}}(\rho_{\text{\tiny SWAP}})$ with the desired confidence of at least $99\%$. The results obtained from both hypothesis-testing protocols are shown in Fig. 6.
@@
Figure 6: Average amount of the fidelity $\mathcal{F}_{\theta=\frac{\pi}{4}}(\rho_{\text{\tiny SWAP}})$ certifiable from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$. For the martingale-based protocol and any given $\mathcal{F}_{0}$ among $\bm{F}_{0}=\{0.5,0.51,\cdots,0.99\}$, we use the $B_{\mathcal{H}}$ determined from Section IV.3 in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{F}_{0}$ from $\bm{F}_{0}$, we solve Section IV.3 by considering the same block size and the level-$2$ outer approximation of $\mathcal{Q}$ introduced in [16]. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{F}_{\theta=\frac{\pi}{4}}(\rho_{\text{\tiny SWAP}})$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{F}_{0}$ in $\bm{F}_{0}$ such that $\mathcal{H}_{\mathcal{F}_{\theta}(\rho_{\text{\tiny SWAP}})\leq\mathcal{F}_{0}}$ is rejected with a $p$-value upper bound less than or equal to $0.01$. Each data point shown in the plot is an average over $30$ such lower bounds. To avoid cluttering the plots, in each line, we show only a small number of markers.
@@
Interestingly, this shows that the martingale-based protocol with the CHSH Bell function of Eq. 4 again performs very well for the self-testing of a Bell state with finite statistics, even though our computation of the corresponding asymptotic confidence-gain rate for $\mathcal{F}_{0}=0.5$ clearly shows that it is suboptimal even for the Bell state, see Fig. 7. What about other partially entangled states? To answer this question, we evaluate the confidence-gain rate derived from both protocols for $\mathcal{F}_{0}\in\bm{F}_{1}=\{\cos^{2}\theta,\cos^{2}\theta+0.9\sin^{2}\theta\}$ with $\theta=\{0^{\circ},1^{\circ},2^{\circ},...,45^{\circ}\}$. This time around, for the martingale-based protocol, we switch to the Bell function of the tilted CHSH Bell inequality of Eq. 19, which is known to facilitate the self-testing of all entangled |$\ket{\psi(\theta)}$. The corresponding results are shown in Fig. 7.
@@
Figure 7: Asymptotic confidence-gain rate $G^{\text{(prot)}}$ based on the family of quantum correlations $\vec{P}_{\theta}$ derived from Section II.2.3, where $\theta=\frac{k\pi}{180}$ rad, $k=\{1,2,\cdots,45\}$. Here, we consider the hypothesis 4, with $\mathcal{F}_{\theta}(\rho_{\text{\tiny SWAP}})=\cos^{2}\theta$, the trivial fidelity achievable without shared entanglement. The gain rate for the martingale-based protocol is computed from Eq. 27 using the CHSH Bell function of Eq. 4 (dashed line, black) and the tilted CHSH Bell function of Eq. 19 (dashed-dotted line, blue) whereas that for the PBR protocol is evaluated from Eq. 36.

### v2

@@ IV.1.2 Dimension certification via negativity certification
As mentioned in Section II.2.1, the correlation $\vec{P}_{\text{CHSH}}$ is insufficient to demonstrate any nontrivial dimension bound. Let us consider, instead, a correlation $\vec{P}_{\text{CGLMP}}$ derived by local measuring the partially entangled two-qutrit state
|$\ket{\Psi}=\frac{1}{\sqrt{2+\zeta^{2}}}(\ket{00}+\zeta\ket{11}+\ket{22}),\quad\zeta=\frac{1}{2}(\sqrt{11}-\sqrt{3}),$ (39a)
@@
where $\varphi^{A}_{x}=\frac{x}{2}$, $\varphi^{B}_{y}=(-1)^{y}\frac{1}{4}$, and {|$\{\ket{j}\}$ is the set of computational basis states. It is known [72, 68] that this strategy gives the maximal CGLMP Bell-inequality violation of $\mathcal{S}_{\text{\tiny CGLMP3}}=1+\sqrt{\frac{11}{3}}\cong 2.91485$. Moreover, the negativity of |$\ket{\Psi}$ can be easily evaluated to give $\cong 0.98358$.
Next, we use the numerically data simulated from $\vec{P}_{\text{CGLMP}}$ to perform a hypothesis testing for Null Hypothesis 1, but now with $\mathcal{N}_{0}\in\bm{N}_{0}=\{0.5,0.51,\cdots,0.98\}$. For the PBR protocol, the computation proceeds in exactly the same way as described above [see the paragraph containing Eq. 38]. However, for the martingale-based protocol, since we do not have an explicit expression like that shown in Eq. 37 for the CGLMP Bell expression, we compute an upper bound on $B_{\mathcal{H}}$ for each given value of $\mathcal{N}_{0}\in\bm{N}_{0}$ according to:
@@
where the CGLMP Bell coefficients $\beta^{ab}_{xy}$ are defined in Eq. 11. Meanwhile, since $P_{xy}=\frac{1}{4}$ and $\beta^{ab}_{xy}\in\{-1,0,1\}$, we again have $\mathfrak{b}_{\pm}=\pm 4$ for $I_{\text{\tiny CGLMP3}}$.
Figure 4: Negativity certifiable from the data observed in a Bell test generating $\vec{P}_{\text{CGLMP}}$. For details on how the plot is generated, see the caption of Fig. 2, but bearing in mind that for the martingale-based protocol, we now use in Eq. 25 the $B_{\mathcal{H}}$ determined from Eq. 40.
@@
and numerically maximize their CGLMP Bell-inequality violation using the heuristic algorithm given in [95]. We denote the corresponding correlation by $\vec{P}_{\tilde{\zeta}}$, compute the corresponding asymptotic confidence-gain rate for both protocols, and plot the results in Fig. 5.
Figure 5: Asymptotic confidence-gain rate $G^{\text{(prot)}}$ based on the quantum correlation $\vec{P}_{\tilde{\zeta}}$ derived from maximizing the CGLMP Bell-inequality-violation of |$\ket{\Psi(\tilde{\zeta})}$, Eq. 41. Here, we consider $\tilde{\zeta}=\{0,0.02,...,1\}$ and the composite hypothesis 1, with $\mathcal{N}_{0}=0.5$. The gain rate for the martingale-based protocol is computed from Eq. 27 using the CGLMP Bell function and the bound $B_{\mathcal{H}}$ determined in Eq. 40 whereas that for the PBR protocol is evaluated from Eq. 36 and setting $\vec{P}_{\mathcal{Q}}$ to $\vec{P}_{\tilde{\zeta}}$. The dashed blue line corresponds to the value $\tilde{\zeta}=\zeta$, cf. Eq. 39a.
Interestingly, even though Fig. 4 suggests that the CGLMP Bell function is very effective in providing a good $p$-value bound against hypothesis 1,  Fig. 5 clearly show that, asymptotically, it is not optimal. The results shown in Fig. 5 further suggest that among the family of two-qutrit states given in Eq. 41, the qutrit signature of |$\ket{\Psi(\zeta)}$, cf. Eq. 39a, could even be the most prominent, when it comes to its DI certification using these hypothesis-testing techniques.
@@ IV.2 Entanglement depth certification
It is easy to verify that $\vec{P}_{\text{GHZ}}$ leads to a violation of the Mermin Bell inequality,  Eqs. 12 and 14, giving the algebraic maximum of $\mathcal{S}_{\text{\tiny Mermin}}=4$. For our simulations, we assume a uniform distribution $P_{xyz}=\frac{1}{4}$ over all measurement settings $x,y,z\in\{0,1\}$ that satisfy $\text{mod}(x+y+z,2)=1$. Then, we test the data against the following composite hypotheses.
Null Hypothesis 2.
$\mathcal{H}_{\text{Sep}}$: In every experimental trial, the underlying state is separable (having an entanglement depth of $1$).
Null Hypothesis 3.
$\mathcal{H}_{\text{2-prod}}$: In every experimental trial, the underlying state is 2-producible, i.e., having an entanglement depth of $2$ or less.
For the martingale-based method, we use Eq. 25 with the Mermin Bell expression of Eq. 12 and the bounds given in Eq. 14, i.e., $B_{\mathcal{H}}=2$ for hypothesis 2 and $B_{\mathcal{H}}=2\sqrt{2}$ for hypothesis 3. Since $\beta^{abc}_{xyz}\in\{-1,0,1\}$, we again have $\mathfrak{b}_{\pm}=\pm 4$. Note that separable states can only generate Bell-local correlations [5], cf. Eq. 2. Thus, for the PBR protocol with hypothesis 2, the optimizing distribution $P^{(k)}_{\star}(abc|xyz)$ for the $k$-iteration can be obtained by solving [cf. Eq. 29]
@@
where $\rho_{i}$, with $i=\{1,2,3\}$ are meant to represent, respectively, the constituent of $\rho$ that is separable with respect to the $A|BC$, $B|AC$, and $C|AB$ bipartitions. In evaluating Eq. 44, we use level $\ell=1$ of the hierarchy introduced in [17]. For both hypotheses, we then evaluate
$R(a,b,c,x,y,z)=\frac{f_{\text{reg}}^{(k)}(abc|xyz)}{P_{\star}(abc|xyz)}$ (45)
for the computation of the test statistic $t_{k}$.
@@
$\mathfrak{P}^{\text{(pbr)}}_{\mathcal{H}_{\text{Sep}}}=0.414958\,N-204.978,\,\,N\in[10^{3},10^{5}],$ (46)
$\mathfrak{P}^{\text{(mart)}}_{\mathcal{H}_{\text{Sep}}}=0.415037\,N,\quad N\in[0,10^{5}]$ (47)
@@
$\mathfrak{P}^{\text{(pbr)}}_{\mathcal{H}_{\text{2-prod.}}}=0.22838\,N-115.22,\,\,N\in[10^{3},10^{5}],$ (48)
$\mathfrak{P}^{\text{(mart)}}_{\mathcal{H}_{\text{2-prod.}}}=0.228447\,N,\quad N\in[0,10^{5}]$ (49)
for the 2-producible hypothesis $\mathcal{H}_{\text{2-prod.}}$ of 3. Consequently, based on this interpolation, even if we only run the Bell test using the strategy of Eq. 42 for 100 trials, there is already sufficient data to certify genuine tripartite-entanglement with a confidence of at least $1-10^{-6}$.
@@ IV.3 Fidelity certification
Null Hypothesis 4.
$\mathcal{H}_{\mathcal{F}_{\theta}(\rho_{\text{\tiny SWAP}})\leq\mathcal{F}_{0}}$: In every experimental trial, the swapped state $\rho_{\text{\tiny SWAP}}$ extractable from the underlying state $\rho$ has a |$\ket{\psi(\theta)}$-fidelity upper bounded by $\mathcal{F}_{0}$, i.e.,
$\mathcal{F}_{\theta}(\rho_{\text{\tiny SWAP}}):=\bra{\psi(\theta)}\rho_{\text{\tiny SWAP}}\ket{\psi(\theta)}\leq\mathcal{F}_{0}.$ (50)
Then, for any given $\theta$ and $\mathcal{F}_{0}\geq\cos\theta$, to apply the PBR protocol, we solve the optimizing distribution $P^{(k)}_{\star}(ab|xy)$ for the $k$-iteration [cf. Eq. 29] via:
@@
where the left-hand side of the last inequality in Eq. 51b consists of some specific linear combination of entries of $\chi_{\ell}[\rho]$, see [87] for details. Then, as with negativity certification, we can evaluate the PBR used in the computation of $t_{k}$ by replacing $f_{\text{reg}}(ab|xy)$ and $P_{\star}(ab|xy)$, respectively, by $f_{\text{reg}}^{(k)}(ab|xy)$ and $P^{(k)}_{\star}(ab|xy)$ in Eq. 30. As for the martingale-based protocol, we first solve
@@
Let us start with the self-testing of a Bell state, corresponding to $\theta=\frac{\pi}{4}$ in Eq. 17a. In this case, we use the CHSH Bell function specified in Eq. 4 and consider $\mathcal{F}_{0}\in\bm{F}_{0}=\{0.5,0.51,\cdots,0.99\}$. For both protocols, by systematically evaluating the $p$-value bounds from the data for each of these $\mathcal{F}_{0}$’s, we determine a lower bound on $\mathcal{F}_{\theta=\frac{\pi}{4}}(\rho_{\text{\tiny SWAP}})$ with the desired confidence of at least $99\%$. The results obtained from both hypothesis-testing protocols are shown in Fig. 6.
@@
Figure 6: Certifiable fidelity $\mathcal{F}_{\theta=\frac{\pi}{4}}(\rho_{\text{\tiny SWAP}})$ from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$. For the martingale-based protocol and any given $\mathcal{F}_{0}$ among $\bm{F}_{0}=\{0.5,0.51,\cdots,0.99\}$, we use the $B_{\mathcal{H}}$ determined from Eq. 52 in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{F}_{0}$ from $\bm{F}_{0}$, we solve Eq. 51 by considering the same block size and the level-$2$ outer approximation of $\mathcal{Q}$ introduced in [17]. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{F}_{\theta=\frac{\pi}{4}}(\rho_{\text{\tiny SWAP}})$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{F}_{0}$ in $\bm{F}_{0}$ such that $\mathcal{H}_{\mathcal{F}_{\theta}(\rho_{\text{\tiny SWAP}})\leq\mathcal{F}_{0}}$ is rejected with a $p$-value upper bound less than or equal to $0.01$.
@@
Interestingly, our results show that the martingale-based protocol with the CHSH Bell function of Eq. 4 again performs very well for the self-testing of a Bell state with finite statistics, even though our computation of the corresponding asymptotic confidence-gain rate for $\mathcal{F}_{0}=0.5$ clearly shows that it is suboptimal even for the Bell state, see Fig. 7. What about other partially entangled states? To answer this question, we evaluate the confidence-gain rate derived from both protocols for [^5] For |$\ket{\psi{(\theta})}$ defined in Eq. 17a, a fidelity of $\cos^{2}\theta$ is always achievable even if Alice and Bob does not share any entanglement; they merely have to prepare |$\ket{00}$ using local operations and classical communication before the Bell test. $\mathcal{F}_{0}=\cos^{2}\theta$ with $\theta=\{0^{\circ},1^{\circ},2^{\circ},...,45^{\circ}\}$. This time around, for the martingale-based protocol, we switch to the Bell function of the tilted CHSH Bell inequality of Eq. 19, which is known to facilitate the self-testing of all entangled |$\ket{\psi(\theta)}$. The corresponding results are shown in Fig. 7.
@@
Figure 7: Asymptotic confidence-gain rate $G^{\text{(prot)}}$ based on the family of quantum correlations $\vec{P}_{\theta}$ derived from Eq. 17, where $\theta=\frac{k\pi}{180}$ rad, $k=\{1,2,\cdots,45\}$. Here, we consider Null Hypothesis 4, with $\mathcal{F}_{\theta}(\rho_{\text{\tiny SWAP}})=\cos^{2}\theta$, the trivial fidelity achievable without shared entanglement. The gain rate for the martingale-based protocol is computed from Eq. 27 using the CHSH Bell function of Eq. 4 (dashed line, black) and the tilted CHSH Bell function of Eq. 19 (dashed-dotted line, blue) whereas that for the PBR protocol is evaluated from Eq. 36.
IV.4 Properties certification via Bell-value certification
The advantage of a fidelity certification based on the SWAP method [89, 87] is that the technique is applicable to a general Bell scenario. However, in the simplest CHSH Bell scenario, it is known that a much tighter lower bound on the Bell-state fidelity can be obtained by considering a more general extraction map. Specifically, Kaniewski showed in [98] that
$\max_{\Lambda_{A},\Lambda_{B}}\min_{\rho_{AB}}\mathcal{F}[(\Lambda_{A}\otimes\Lambda_{B})(\rho_{AB}),\ket{\psi_{\text{MES}}}$ $\bra{\psi_{\text{MES}}}])\geq$ (53)
$\frac{1}{2}+\frac{1}{2}\frac{\mathcal{S}_{\text{\tiny CHSH}}-\beta^{*}}{2\sqrt{2}-\beta^{*}},$
where $\Lambda_{A}$, $\Lambda_{B}$ are local extraction maps acting, respectively, on Alice’s and Bob’s subsystem, while $\beta^{*}:=\frac{16+14\sqrt{2}}{17}\approx 2.1058$ is the threshold CHSH value for which the fidelity bound becomes trivial.
To take advantage of Eq. 53, we can first perform a hypothesis testing based on the following null hypothesis.
Null Hypothesis 5.
$\mathcal{H}_{\mathcal{S}_{\text{\tiny CHSH}}(\rho)\leq\mathcal{S}_{0}}$: In every experimental trial, the underlying state and measurements give a CHSH value $\mathcal{S}_{\text{\tiny CHSH}}$ less than or equal to $\mathcal{S}_{0}$.
Specifically, using the same set of data generated for the analysis in Section IV.1.1 and Section IV.3, we perform composite hypothesis testing for Null Hypothesis 5 with $\mathcal{S}_{0}\in\{2,2+\Delta\mathcal{S},2+2\Delta\mathcal{S},\cdots,2\sqrt{2}-\Delta S\}$, where $\Delta\mathcal{S}=\frac{2(\sqrt{2}-1)}{50}$.
In particular, for the martingale-based protocol, we can simply use Eq. 25 with $\mathfrak{b}_{\pm}=\pm 4$ and $B_{\mathcal{H}}=\mathcal{S}_{0}$. Could one also employ the PBR protocol, which does not usually presuppose any Bell-like inequality, for the current hypothesis testing? This is indeed possible. To this end, one may solve the optimizing distribution $P^{(k)}_{\star}(ab|xy)$ for the $k$-iteration [cf. Eq. 29] of the PBR protocol via:
s.t. $\sum_{a,b,x,y}(-1)^{xy+a+b}P(ab|xy)\leq\mathcal{S}_{0},$ (54b)
with or without imposing the SDP constraints of Eq. 9c. The results obtained from these tests are shown in Fig. 8.
Figure 8: Certifiable Bell-CHSH violation $\mathcal{S}_{\text{\tiny CHSH}}$ from the data observed in a Bell test generating $\vec{P}_{\text{CHSH}}$. For the martingale-based protocol and any given $\mathcal{S}_{0}$ among $\bm{S}_{0}=\{2+k\Delta\mathcal{S}\}_{k=0}^{49}$, we use Eq. 54b in Eq. 25 to upper-bound $p^{\text{(mart)}}$ after every block of $N_{\text{blk}}=500$ trials, thereby generating $200\times 50$ upper bounds on $p^{\text{(mart)}}$ for a complete Bell test. For the PBR protocol and a given $\mathcal{S}_{0}$ from $\bm{S}_{0}$, we solve Eq. 54 by considering the same block size. Then, we obtain $199\times 50$ upper bounds on $p^{\text{(pbr)}}$ from Eqs. 34, 35 and 32. To determine the lower bound on the underlying $\mathcal{S}_{\text{\tiny CHSH}}$ with the desired confidence of $\gamma\geq 99\%$, we look for the largest $\mathcal{S}_{0}$ in $\bm{S}_{0}$ such that $\mathcal{H}_{\mathcal{S}_{\text{\tiny CHSH}}\leq\mathcal{S}_{0}}$ is rejected with a $p$-value bound being less than or equal to $0.01$. A separate calculation shows that if we impose Eq. 9c in addition to Eq. 54b, one may find a visually indistinguishable difference ($<5\times 10^{-4}$) in the certifiable $\mathcal{S}_{\text{\tiny CHSH}}$.
Using each lower bound on $\mathcal{S}_{\text{\tiny CHSH}}$ certified from the data, Eq. 53 immediately translates to a lower bound on the Bell-state fidelity with the desired confidence. For a direct comparison with the efficacy of the SWAP-based approach adopted in Section IV.3, we plot in Fig. 9 the Bell-state fidelity certifiable using the two approaches. As expected, the tighter Bell-state fidelity lower bound provided by Eq. 53 also facilitates a considerably tighter lower bound when one has access to only a finite amount of data.
Figure 9: Comparison of the Bell-state fidelity certifiable via the SWAP-based approach presented in Section IV.3 and that via Eq. 54 and $\mathcal{S}_{\text{\tiny CHSH}}$-value certification.
It is also worth noting that in computing these PBR bounds, the computation may be further simplified by regularizing the relative frequency $\vec{f}_{\text{reg}}$ using only the nonsignaling constraint of Eq. 7, instead of the quantum approximation via Eq. 9c. For the lower bounds on $\mathcal{S}_{\text{\tiny CHSH}}$ presented in Fig. 8, this further simplification was found to give, unsurprisingly, a worse lower bound, but with a deviation bounded by $8\times 10^{-3}$. Of course, the lower bounds on $\mathcal{S}_{\text{\tiny CHSH}}$ can also be used to bound other desired properties. For example, Fig. 2 can equivalently be obtained by combining Eq. 10 with the results shown in Fig. 8.